OpenStudy (anonymous):

If (x)/(2y) = 3, then what is the value of x^2- 36y^2

OpenStudy (zzr0ck3r):

this is all the information they give?

OpenStudy (zzr0ck3r):

x^2 + -36(x/6)^2= x^2-36*(x^2)/36 = x^2-x^2 = 0

OpenStudy (anonymous):

ok thanks